# Inner crust of neutron stars: Unified description of structure and superconductivity

### v1

@@ IV.1 Characteristic physical lengths of protons
In Eq. (14), the proton Fermi velocity is $v_{Fp}=\hbar k_{p}/m_{p}$, the Euler constant is $\ln\gamma=0.577$ and $\Delta_{p}=\Delta_{p}(n,x)$ is the superconducting energy gap. The proton Fermi wavenumber is given by
@@
where the quantity $Y_{p}$ (the proton fraction averaged over the WS cell) utilized in LimHolt2021 is given in variables of the present paper as
@@
[…]
A thermodynamic condition for the existence of the coherent (Cooper) pairing is given by Eq. (15). In order to reveal what part of the 1N phase is expected to be superconducting, we plot in Fig. 7 the superconducting gap energy as function of the mean baryon density across the entire density range of 1N phase.
Figure 7: The superconducting energy gap $\Delta_{p}$ adopted from Fig. 7 of LimHolt2021 , where the $k_{Fp}$-dependence has been converted into the $n_{b}$-dependence with the help of the functions shown in Figs. 1–4 and Eq. (16).
Figure 7 shows that at typical $k_{B}T\sim 0.01$ MeV, the thermodynamic condition given in Eq. (15) does hold in the entire 1N phase and therefore this phase is expected to host superconducting protons.
The basic magnetic properties of superconducting matter in the 1N phase can be characterized by the ratio $d_{L}/\xi$ between the interlayer spacing $d_{L}$, where
@@
defines the square of the London penetration depth of the magnetic field for the bulk matter. By using in Eq. (19) the value of the superconducting proton density averaged over the WS cell, we obtain the lower bound on the magnetic penetration depth. If neutrons are introduced between the slabs (such as in the realistic configuration of the 1N phase), the magnetic penetration depth might increase because in this case the superconductor becomes intrinsically multiply-connected as show panels (a) and (b) in Fig. 8.
Figure 8: Schematic representation of a unit length of the ideally ordered 1N phase, which currently is not excluded as a real state of the inner crust at the relevant density range of the inner crust of neutron stars. The three dots in panels (a) and (b) represent the ideal lattice of slabs. Panel (a) shows a configuration with the magnetic field parallel to the slabs. Panel (b) displays a configuration with the magnetic field parallel to the slabs with dashed lines representing the normal core of the magnetic flux tube in the proton superconductor. Panel (c) depicts a single slab within the magnetic induction with arrows representing the neutron-proton momentum lag vector at equilibrium.
The magnetic flux through the 1N phase is nonquantized in the case shown in Fig. 8 (a) and is quantized in Fig. 8 (b). Notice that the anisotropy of the 1N phase induces a direction-dependence of the London penetration depth, in contrast to the isotropic Eq. (19). As has been pointed out in Kobyakov2018 , for ideally ordered 1N phase without tunneling of Cooper pairs between adjacent superconducting slabs in a parallel magnetic field as shown in Fig. 8 (a), the magnetic penetration depth is macroscopically large.
The GL parameter $\lambda/\xi$ indicates whether the superconductor is type-I or type-II Kobyakov2020 . The numerical results reported in this paper allow to evaluate the GL parameter along with other characteristic quantities of the proton superconductor in the 1-dimensional pasta phase.
Figure 9 shows the dependence of the parameters $d_{L}$, $r_{N}$ and $\xi$ in the 1N phase on the mean baryon density $n_{b}$, where we also plot the scaled London penetration depth $\lambda$, the characteristic screening length $k_{TFe}^{-1}$ in the electron background and the inverse Fermi wavenumber of the electrons $k_{e}^{-1}$. Here,
@@
with $n_{e}$ and $\mu_{e}$ given by Eqs. (72) and (87), correspondingly, so ${\partial n_{e}}/{\partial\mu_{e}}=k_{e}^{2}/\pi^{2}\hbar c$, where
@@
In Fig. 9, we observe that $\lambda/\xi>\sqrt{2}$ in the entire range of $n_{b}$ corresponding to 1N phase and therefore the 1N phase is expected to be in the type-II regime of superconductivity.
@@
Figure 9: The characteristic parameters of proton superconductor in 1N phase calculated from Figs. 1-4 and supplemented by the data on $\Delta_{p}(k_{Fp})$ from Fig. 7 of LimHolt2021 . The distance between the nearest interfaces of two adjacent slabs is $d_{L}$, $\lambda$ is the lower bound on the London penetration depth of the magnetic field into the superconducting protons, $\xi$ is the coherence length of protons at temperatures well below $k_{B}T$, $2r_{N}$ is the width of the slab (also given in Fig. 2), $k_{TFe}^{-1}$ is the characteristic screening length in the electron background and $k_{e}^{-1}$ is the electron Fermi wavenumber.
Physically, $\xi$ provides an estimate for the size of the normal core of the quantum magnetic flux tube. In the normal core, the superconducting density smoothly goes to zero. In the nuclear system, the normal cores are filled by the normal protons, so the sum of the superconducting and the normal proton densities is constant. There is yet another physical interpretation of $\xi$ as characteristic size of the Cooper pair. General considerations suggest that the slab can be thought of a three-dimensional superconducting domain when size of the Cooper pair $\xi$ is smaller than the domain width along the smallest dimension, which is $2r_{N}$ in case of 1N phase. When the Cooper pair $\xi$ is larger than $2r_{N}$, the slab represents effectively two-dimensional system.
Another important aspect of the problem is whether the Cooper pair size $\xi$ is smaller or larger than the spacing $d_{L}$ between two adjacent slabs. If two slabs are a pair of parallel two-dimensional nuclear films separated by a distance $d_{L}$ smaller than the size $\xi$ of the Cooper pair, $d_{L}<\xi$, then it is natural to expect that the phase fields of the order parameter corresponding to the two slabs are synchronized i.e. mutually coherent. On the contrary, when $d_{L}>\xi$, it is natural to expect that the slabs are mutually incoherent, i.e. their phase fields are independent because they are separated by a locally low-density region, which contains large superconducting phase fluctuations that destroy the coherence. Therefore the results shown in Fig. 9 suggest that for $n_{b}\gtrsim 0.48n_{0}$ the slabs are coupled by the Josephson effect because $\xi>d_{L}$.
It is seen that the individual slabs have two-dimensional character for $n_{b}\lesssim 0.54n_{0}$, while for $n_{b}\gtrsim 0.54n_{0}$ the slabs are effectively three-dimensional. We observe that the approach used by Zhang and Pethick in ZhangPethick2021 is valid in the range $0.48n_{0}\lesssim n_{b}\lesssim 0.54n_{0}$.
In the range $n_{b}\lesssim 0.48n_{0}$ the slabs are not significantly coupled by the Josephson effect because $\xi<d_{L}$. Thus, the approach used by Kobyakov in Kobyakov2018 is valid in the range $0.37n_{0}\lesssim n_{b}\lesssim 0.48n_{0}$.
IV.2 Natural stress of the superconducting-superfluid crust in neutron stars
In the standard astrophysical picture, neutron stars observable as pulsars are characterized by the electromagnetic emission related to rotation of the dipolar magnetic field in space because the latter is misaligned with the spinning axis. As a result, rotation of the conglomerate of charged particles inside pulsar continuously slows down due to the emission losses. This is not the case for the superfluid neutrons, which are expected to slow down via abrupt jumps of the rotation frequency due to the quantum character of the rotational energy of neutrons, namely the pinning of the superfluid neutron vortices prevents their continuous outflow and the total number of vortices changes abruptly (as a vorticity discharge) when avalanche of vortices suddenly unpins. Thus, the mutual rotational evolution of the nonsuperfluid and superfluid particles naturally includes situations when there is an adiabatic perturbation of the macroscopic flow of neutrons in the rest frame of reference of the center-of-mass of the slab. A relevant question in this case is, whether the perturbation of the neutron superflow in the magnetic field $\mathbf{H}$ would lead to a magnetohydrodynamic stress?
@@ IV.2 Natural stress of the superconducting-superfluid crust in neutron stars
Stresses of the superconducting-superfluid crust in neutron stars can be conveniently studied in the framework of the stress tensor approach. Tensor of the momentum and energy, or briefly, the stress tensor, contains information about hydrodynamic and magnetic content in the system. It is useful for studies of pressure forces and the interaction of the fluid with the magnetic field. Often, it is convenient to study the stress tensor for some specific form of motion and specific initial conditions for the magnetic field. For instance, the case that can be loosely described as magnetostatic-hydrostatic was studied by Easson and Pethick in EassonPethick1977 , while the magnetostatic-hydrodynamic situation was studied by Sedrakian and Sedrakian in Sedrakian1995 in the context of the uniform nuclear matter in the core of neutron stars and later by Kobyakov in Kobyakov2018 in the context of nonuniform matter in the inner crust of neutron stars.
The pioneering work EassonPethick1977 was done for the case of a uniform isotropic superconductor threaded by the quantized magnetic flux lines, while there is no macroscopic flow of matter. In the setting utilized in EassonPethick1977 , a virtual change in the parallelepiped leads to a perturbation of the total density of matter. Thus, in uniform neutron star matter with nucleon densities denoted by $n_{p}=xn$ and $n_{n}=(1-x)n$, the magnetic part of the stress tensor (see Eq. (8) in EassonPethick1977 ) reads
$\sigma^{\rm mag,\,unif}_{ij}$
$=\left[F^{\rm mag}-n\left(x\frac{\partial}{\partial n_{p}}+(1-x)\frac{\partial}{\partial n_{n}}\right)F^{\rm mag}-\frac{1}{4\pi}\mathbf{H}\cdot\mathbf{B}\right]\delta_{ij}$
$+\frac{1}{4\pi}H_{i}B_{j},$ (22)
where $F^{\rm mag}$ is the magnetic part of the free energy (see Eq. (5) in EassonPethick1977 ) defined as the piece of the total free energy $F$ that is nonzero only when the magnetic induction is present:
$F=F^{\rm matter}+F^{\rm mag}(\mathbf{B}),$ (23)
and where the dependence on the matter density is implicit. Equation (22) holds independently of whether the protons are superconducting or not. The information regarding the superconductivity and the coupled superfluidity is contained inside $F^{\rm mag}$.
In the following work, scenarios involving some form of a macroscopic fluid flow were studied. The approach of Sedrakian and Sedrakian Sedrakian1995 focused on interaction of the superconducting protons with the superfluid rotation of neutrons, which generates the magnetization due to the entrainment effect.
The intrinsic momentum coupling between superfluid neutrons and superconductor protons is a phenomenon known as the entrainment in the superconducting-superfluid uniform nuclear matter. The entrainment contribution to the total energy density depends on the electromagnetic vector potential $\mathbf{A}$ through the gauge-invariant difference of the gradient of the phase of the proton superconducting order parameter $\nabla\phi_{p}$ and $\mathbf{A}$. This contribution may be understood as the one coming from the $F^{\rm mag}(\mathbf{B})$ dependence.
As a next step, recently, effects of a macroscopic flow of superfluid neutrons in nonuniform matter of the inner crust of neutron stars were studied by Kobyakov Kobyakov2018 . Notice that the elastic properties are irrelevant for the aims of finding the force due to the combined action of the magnetic field and the neutron superflow in Kobyakov2018 .
A natural feature of the inner crust matter is that the superconductor though it is still type-II, is extremely anisotropic. As it follows from our calculations, this happens at the mean baryon number density close to the tip of 1N phase, where $\xi<d_{L}$, see Fig. 9. This is because in 1N phase the proton supercurrents are possible only inside the nuclear slabs of the ideally ordered 1N phase and not between them via the pure neutron liquid.
In case when $\xi<d_{L}$, depending on the relative orientation of the structure and the applied magnetic field, the magnetic flux penetrates either in the form of a quantized flux tube array, as in the case shown in Fig. 8 (b), or in the form of a nonquantized magnetic flux through the voids between the slabs, as in the case shown in Fig. 8 (a). Notice that the magnetic induction is almost equal to the magnetic field inside the slab, because the condition
@@
does hold for the entire domain of the 1N phase, as shows Fig. 8.
Let us turn to the special (but natural) case, when the magnetic field is parallel to the slab surface, as shows Fig. 8 (a). In this case, ignoring the anisotropy of the elastic energy and utilizing the notation used for description of the pasta phases, we find:
[…]
$=\left[F^{\rm mag}(u,n,x,\mathbf{A})-n_{b}\frac{\partial F^{\rm mag}(u,n,x,\mathbf{A})}{\partial n_{b}}\right.$
$\left.-\frac{1}{4\pi}\mathbf{H}\cdot\mathbf{B}\right]\delta_{ij}+\frac{1}{4\pi}H_{i}B_{j},$
where the entrainment contribution $-\frac{1}{2}mn_{np}\mathbf{w}^{2}\delta_{ij}$ (introduced in Eq. (8) of Kobyakov2018 ) is contained in $F^{\rm mag}(u,n,x,\mathbf{A})$. The momentum lag is defined as
$\mathbf{w}=\frac{1}{m}\left[\left(\mathbf{p}_{p}-\frac{e}{c}\mathbf{A}\right)-\mathbf{p}_{n}\right],$ (26)
where $\mathbf{p}_{q}=\hbar\nabla\phi_{q}$, ($q=p,n$), with $\phi_{q}$ being the superfluid phase of the order parameters of the nucleons of type $q$. The entrainment contribution gives rise to a characteristic force operating when the neutron-proton lag $\mathbf{w}$ is nonzero Kobyakov2018 . A brief account of the derivation of this force from the basic energy functional is given in Appendix E, where the magnetohydrodynamic stress tensor is denoted as $\Pi_{ij}$. The expression for the force was given in Eq. (40) of Kobyakov2018 (see also a corrected expression in Kobyakov2023 ).

### v2

@@ VI.2 Characteristic physical lengths of protons
In Eq. (18), the proton Fermi velocity is $v_{Fp}=\hbar k_{p}/m_{p}$, the Euler constant is $\ln\gamma=0.577$ and $\Delta_{p}=\Delta_{p}(n,x)$ is the superconducting energy gap. The proton Fermi wavenumber reads
@@
where the quantity $Y_{p}$ (the proton fraction averaged over the Wigner-Seitz cell) utilized in LimHolt2021 is given in variables of the present paper as
@@
In order to reveal what part of lasagna is expected to be superconducting, we plot in Fig. 7 the superconducting gap energy as function of the mean baryon density across the entire density range of lasagna.
Figure 7: The superconducting energy gap $\Delta_{p}$ in lasagna, adopted from Fig. 7 of LimHolt2021 , where the $k_{Fp}$-dependence has been converted into the $n_{b}$-dependence with the help of the functions shown in Figs. 1–4 and Eq. (20).
Figure 7 shows that at typical $k_{B}T\sim 0.01$ MeV, the thermodynamic condition given in Eq. (19) does hold in the entire domain of lasagna.
The basic magnetic properties of superconducting matter in lasagna can be characterized by the ratio $d_{L}/\xi$ between the interlayer spacing $d_{L}$, where
@@
defines the square of the London penetration depth of the magnetic field for the bulk matter. By using in Eq. (23) the value of the superconducting proton density averaged over the Wigner-Seitz cell, we obtain the lower bound on the magnetic penetration depth. If there are neutrons between the slabs, the magnetic penetration depth might increase because in this case the superconductor becomes topologically different from uniform superconductor, as show panels (a) and (b) in Fig. 8.
Figure 8: Schematic representation of a unit length of ideally ordered lasagna, which currently is not excluded as a real state of the inner crust. Three dots in panels (a) and (b) express the presence of the ideal lattice of slabs. Panel (a) shows a configuration with the magnetic field parallel to the slabs. Panel (b) displays a configuration with the magnetic field parallel to the slabs with dashed lines representing the normal core of the magnetic flux tube in the proton superconductor. Panel (c) depicts a single slab within the magnetic induction with arrows representing the neutron-proton momentum lag vector at equilibrium.
The magnetic flux through lasagna is nonquantized in the case shown in Fig. 8 (a) and is quantized in Fig. 8 (b). Notice that the anisotropy of lasagna induces a direction-dependence of the London penetration depth, in contrast to the isotropic Eq. (23).
Figure 9 shows the dependence of the parameters $d_{L}$, $r_{N}$ and $\xi$ in lasagna on the mean baryon density $n_{b}$, where we also plot the scaled London penetration depth $\lambda$, the characteristic screening length $k_{TFe}^{-1}$ in the electron background and the inverse Fermi wavenumber of the electrons $k_{e}^{-1}$. Here,
@@
with $n_{e}$ and $\mu_{e}$ given in Eqs. (78) and (91), correspondingly, so ${\partial n_{e}}/{\partial\mu_{e}}=k_{e}^{2}/\pi^{2}\hbar c$, where
@@
In Fig. 9, we observe that $\lambda/\xi>\sqrt{2}$ in the entire range of $n_{b}$ corresponding to lasagna and therefore lasagna is expected to be in the type-II regime of superconductivity.
@@
Figure 9: The characteristic parameters of proton superconductor in lasagna calculated from Figs. 1-4 and supplemented by the data on $\Delta_{p}(k_{Fp})$ from Fig. 7 of LimHolt2021 . The distance between the nearest interfaces of two adjacent slabs is $d_{L}$, $\lambda$ is the lower bound on the London penetration depth of the magnetic field into the superconducting protons, $\xi$ is the coherence length of protons at temperatures well below $\Delta_{p}/k_{B}$, $2r_{N}$ is the width of the slab (also given in Fig. 2), $k_{TFe}^{-1}$ is the characteristic screening length in the electron background and $k_{e}^{-1}$ is the electron Fermi wavenumber.
It is seen in Fig. 9, that the individual slabs have 2-dimensional character for $0.37n_{0}\lesssim n_{b}\lesssim 0.54n_{0}$ (when $\xi>2r_{N})$, while for $0.54n_{0}\lesssim n_{b}\lesssim 0.56n_{0}$ the slabs are effectively 3-dimensional (when $\xi<2r_{N}$). The information about dimensional character of the slabs is essential for calculations involving the magnetic vortex structure and the critical magnetic field.
Another aspect of the problem is the interaction between the neighboring slabs, which can be quantified by the ratio of $\xi$ to $d_{L}$. In the range of $0.37n_{0}\lesssim n_{b}\lesssim 0.48n_{0}$ (when $\xi<d_{L}$), the slabs are not significantly coupled by the Josephson effect and therefore, the appropriate picture is the layered discreet superconductivity. However, at higher densities, in the range of $0.48n_{0}\lesssim n_{b}\lesssim 0.56n_{0}$ (when $\xi>d_{L}$), the appropriate picture is the anisotropic 3-dimensional superconductivity.
In a recent work by Zhang and Pethick ZhangPethick2021 , it was assumed that the individual slabs have 2-dimensional character and that the neighboring slabs are coupled by the Josephson effect and, thus, the 3-dimensional anisotropic picture of superconductivity was used. Figure 9 shows that the approach used in ZhangPethick2021 is validated by our numerical results in the range of $0.48n_{0}\lesssim n_{b}\lesssim 0.54n_{0}$. In contrast, in an earlier work by Kobyakov Kobyakov2018 , it was assumed that the neighboring slabs are not coupled by the Josephson effect and thus, the discreet layered picture of superconductivity was used. This approach is validated in the range of $0.37n_{0}\lesssim n_{b}\lesssim 0.48n_{0}$.
VII Astrophysical implications
VII.1 Overview of calculation of stress in superconducting matter of neutron stars
@@
Tensor of the momentum and energy, or briefly, the stress tensor, contains information about hydrodynamic and magnetic content in the system. It is useful for studies of pressure forces and of interaction of the fluid with the magnetic field. Often, it is convenient to study the stress tensor for some specific form of motion and specific initial conditions for the magnetic field. For instance, the case that can be loosely described as magnetostatic-hydrostatic was studied by Easson and Pethick in EassonPethick1977 , while the magnetostatic-hydrodynamic situation was studied by Sedrakian and Sedrakian in Sedrakian1995 in the context of the uniform nuclear matter in the core of neutron stars and later by Kobyakov in Kobyakov2018 in the context of nonuniform matter in the inner crust of neutron stars.
The pioneering work EassonPethick1977 was done for the case of a uniform isotropic superconductor threaded by the quantized magnetic flux lines, while there is no macroscopic flow of matter. In the following work, scenarios involving some form of a macroscopic fluid flow were studied. The approach of Sedrakian and Sedrakian Sedrakian1995 focused on interaction of the superconducting protons with the superfluid rotation of neutrons, which generates the magnetization due to the entrainment effect. Intrinsic momentum coupling between superfluid neutrons and superconducting protons is a phenomenon known as the entrainment in the superconducting-superfluid uniform nuclear matter. The entrainment contribution to the total energy density depends on the electromagnetic vector potential $\mathbf{A}$ through the gauge-invariant difference of $\mathbf{A}$ and the gradient of the phase of the proton superconducting order parameter $\nabla\phi_{p}$. As a next step, Kobyakov Kobyakov2018 studied some effects of a macroscopic flow of superfluid neutrons in lasagna. Notice that the magnetic induction is almost equal to the magnetic field inside the slab, because, as shows Fig. 9, the condition
@@
holds for the entire domain of lasagna.
The mutual rotational evolution of the nonsuperfluid and superfluid particles naturally includes situations when there is an adiabatic perturbation of the macroscopic flow of neutrons in the rest frame of reference of the center-of-mass of the slab. A relevant question in this case is, whether the perturbation of the neutron superflow in the magnetic field $\mathbf{H}$ would lead to a magnetohydrodynamic stress?
This question has been raised in the earlier work Kobyakov2018 . The basic idea in Kobyakov2018 is that the system is a toroidal volume containing ideal lasagna. The slabs are oriented along the field lines of the toroidal component $\mathbf{B}$ of the stellar magnetic induction, as schematically shown in Fig. 10 (a).
Figure 10: (Color online). Panel (a): A schematic representation (not to the scale) of the magnetic-equatorial cross-section of the star, highlighting lasagna located between the outer crust (when mean baryon density $n_{b}$ is smaller than the neutron drip density) and the core (when $n_{b}>n_{0}\eta_{\rm uni}^{*}$) and threaded by the toroidal lines of the stellar magnetic induction $\mathbf{B}$. Thin solid rectangle denoted as $\gamma$ represents the cut piece of the toroidal volume shown in panel (a) of Fig. 8. Panel (b): Enlarged view of 2-dimensional projection of 3-dimensional piece of volume $\gamma$ “cut” from panel (a), where $D$ is the height of the column discussed in Eq. (45). The planar nature of the system seen in panel (b) and in Fig. 8 is due to the very large curvature radius of the structure as compared with the structure lattice period. The periodic boundary conditions in the plane parallel to the slabs, which are used in our calculations are essential for existence of the neutron superflow. The parameter $L$ was introduced in Eq. (62).
Let us now “cut” a piece of lasagna, with spatial dimensions $D\times L^{2}$, as schematically shows thin rectangle in Fig. 10 (a). This rectangle is enlarged in Fig. 10 (b) in order to clearly define the physical system under consideration. By cutting a piece with $2r_{N}\ll L\ll R_{*}$, where $R_{*}$ is the stellar radius, we convert the physical problem of the inner crust shown in Fig. 10 (a) into an idealized model where the slabs are strictly flat as shown in Fig. 10 (b). This situation corresponds to the model shown in Fig. 8 (a). We assume periodic boundary conditions for the nucleon number currents in the plane parallel to the slab, so that the neutron and proton flows parallel to the slab surface are possible. If the condition $d_{L}>\xi$ is satisfied, the slabs in Fig. 8 (a) are independent. We assume this condition holds and therefore it is sufficient to consider only a single slab, as shown in Fig. 8 (c).
Following Kobyakov2018 , we assume that the temperature is low enough, so that the superfluid nucleon densities are equal to the bulk nucleon densities $n_{\alpha}$, $(\alpha=p,n)$. For convenience, let us remind that the superconducting and superfluid S-wave order parameters, which can be generally written as $\sqrt{n_{\alpha}}\exp{\mathrm{i}\phi_{\alpha}}$, consist of the superfluid density $n_{\alpha}$ and the superfluid phase $\phi_{\alpha}$. The fluid specific momenta are $\mathbf{p}_{\alpha}=\hbar\nabla\phi_{\alpha}$. The lag of superfluid momenta scaled by the particle mass, is defined as
$\mathbf{w}=\frac{1}{m}\left[\left(\mathbf{p}_{p}-\frac{e}{c}\mathbf{A}\right)-\mathbf{p}_{n}\right].$ (27)
The energy density of the system shown in Fig. 8 (c) reads
$w_{\rm tot}^{\rm matt}=w^{\rm nuc}+w_{\rm kin}^{\rm nuc}+w_{\rm Coul}^{\rm ep}+w^{\rm e}+w^{\rm no},$ (28)
Here, $\boldsymbol{\pi}_{p}=\mathbf{p}_{p}-e\mathbf{A}/c$ is the gauge-invariant proton momentum, $\mathbf{A}$ is the electromagnetic vector potential and $m=m_{p}\approx m_{n}$ and $\mathbf{B}=\nabla\times\mathbf{A}$.
The equations of motion generated from the energy functional with density given in Eq. (28) have the form of the continuity equation and the Newtown’s second law for each of the nucleon species Kobyakov2018 :
$\partial_{t}n_{\alpha}+\nabla\cdot\mathbf{J}_{\alpha}=0,$ (30)
where $\mathbf{J}_{\alpha}$ are the number currents of each species of nucleons. Combining the momentum equations and using the incompressibility condition $\nabla\cdot\mathbf{J}_{\alpha}=0$, one arrives at the expression for the stress tensor.
Using the geometry shown in Fig. 8 (c), the equilibrium vector potential is $\mathbf{A}_{0}=-\mathbf{\hat{y}}B_{0}z$ and the unperturbed superfluid gauge-invariant momentum lag $\mathbf{w}_{0}$ reads
$\mathbf{w}_{0}=(e/mc)\mathbf{\hat{y}}B_{0}z,$ (32)
see equation (35) in Kobyakov2018 . The total force between the magnetic field and the slab is zero in equilibrium. This also can be seen from Eq. (41) with $\mathbf{w}=\mathbf{w}_{0}$, or, $\delta\mathbf{w}=0$.
VII.2 Calculation of stress in a symmetric setting
In the remainder of this section we will study perturbation of the force as a result of perturbation of the neutron superflow. The expression for the force was given in Eq. (40) of Kobyakov2018 (see also a corrected expression in Kobyakov2023 ). As compared with the work done in Kobyakov2018 , here we will specify the quantitative evaluation of the force based on the microscopic parameters calculated in Sec. IV and generalize the expression for the force for the case when $\mathbf{A}_{0}$ and perturbation $\delta\mathbf{w}$ are not parallel. Following Kobyakov2018 and in the coordinates shown in Fig. 8 (c), the relevant force with $L=1$ cm reads
where the stress tensor is
The hydrodynamic pressure reads
$P=p-\rho_{np}^{*}\mathbf{w}^{2}/2,$ (35)
where $p$ is the pressure without the contributions due to the matter flows, as given in equation (32) in Kobyakov2018 . $P$ is generally defined as the material contribution to $\Pi_{ik}$ associated with the isotropic tensor $\delta_{ik}$. Here, $\rho_{np}^{*}=mn_{np}-mn_{p}\theta_{p}-mn_{n}\theta_{n}$ is the entrainment mass density (in the Skyrme model, $\theta_{p}=\theta_{n}=0$), $n_{np}$ is the entrainment number density and $P$ is the total physical pressure (the hydrodynamic pressure). We evaluate the entrainment number density from (see Eq. (49) in KobyakovEtAl2017 )
$n_{np}=\frac{2}{n_{0}}\left(1-\frac{m}{m_{0}^{*}}\right)n_{n}n_{p},$ (36)
where $n_{p}$ and $n_{n}$ are the nucleon number densities in uniform matter, $m_{0}^{*}$ is so-called nucleon isovector effective mass, which reads (see Eq. (3.22) in ChabanatEtal1997 )
$m_{0}^{*}=\frac{m}{1+({m}/{4\hbar^{2}})n_{0}\left[t_{1}\left(x_{1}+2\right)+t_{2}\left(x_{2}+2\right)\right]}.$ (37)
${m_{0}^{*}}=\left.\tilde{m_{q}}\right|_{n_{q}=0}.$ (38)
It remains to evaluate the total force as resultant of the forces applied to each of the slabs in the lattice (see Fig. 8 (a)). From independence of the slabs, the total force acting on a 1 ${\rm cm}^{2}\times D$ column of lasagna is a sum of (integer part of) ${D}/{2r_{c}}$ equal terms, or
where $D$ is the height of the column measured along the normal direction of the slabs, as shows panel (b) in Fig. 10.